Determine whether the following is an expression or an equation.
-3x+2-1=4x

Found 3 solutions by Alan3354, amfagge92, ikleyn:
Answer by Alan3354(69443)

It has an equal sign --> an equation.

Answer by amfagge92(93)

EQUATION
X=1/7
CHECK=-3x+1-4x=0
-3(1/7)+1-4(1/7)=0
-3+7-4=0
THANKS ALL.

Answer by ikleyn(52778)

.
an equation, without doubt.